OpenStudy (anonymous):

what is the quotientof 4 and 2/5

OpenStudy (anonymous):

To make multiplying the two of the fractions easier lets out 4 over one. Thats make 4 a fraction but it still equals 4. To divide fractions you take 4/1 and multiply it by the reciprocal of 2/5 which is 5/2.

OpenStudy (anonymous):

I meant to make dividing fractions easier

OpenStudy (shane_b):

@lkjhgfds: Is the question supposed to be:\[\frac{4}{\frac{2}{5}}=?\]or \[4\frac{2}{5}=?\]